27.Choose the aqueous solution below with the lowest freezing point. These are all solutions of nonvolatile solutes and you should assume ideal vant Hoff factors where applicable. A. 0.075 m NaI B. 0.075 m (NH4)3PO4 C. 0.075 m NaBrO4 D. 0.075 m LiCN E. 0.075 m KNO2
Answer #1

where, ??_ the amount the freezing temperature decreases the freezing point depression constant which depends on the solvent (for water, Kf =1.86°C/m) m- the molality of the solution ?the number of particles formed when that compound dissolves (for covalent compounds, this number is always 1) Lets calculate depression in freezing point of all solution, solution having largest ?? will have lowest freezing point: for all solutions m = 0.075 m A) Nal: i - 2 because it dissociated into two ions, Na+ and I- Tf -1.86?0.0752-0.279 ? B) (NH4)3P04: i = 4 (3NH4+ and 1P04-ions) ?Tf -1.86*0.075*4-0 .558 oC C) NaBrO4: i 2 (1Na+ and 1 Br04- ions) Tr = 1.86 0.075°2 = 0.279 C D) L?CN: i = 2 (Li+ and-CN) ?Tf -1.860.075 2-0.279 °C E) KNO2: i-2, (K+ and NO2- ions) Tr = 1.860.075°2= 0.279°C Thus largest ??? is in case of (NH4)3P04 hence the solution B) will have lowest freezing point. Answer is option B.